Six shirts and two ties are purchased for $88.68. Then, four of the same-priced shirts and three of the same-priced ties are bought for $68.27. What are the prices of a single shirt and tie?
Answer by mananth(16946)

each shirt x $
each tie y $

6 x + 2 y = 88.68 .............1
4 x + 3 y = 68.27 .............2
Eliminate y
multiply (1)by -3
Multiply (2) by 2
-18 x -6 y = -266.04
8 x 6 y = 136.54
Add the two equations
-10 x = -129.5
/ -10
x = 12.95
plug value of x in (1)
6 x + 2 y = 88.68
77.7 + 2 y = 88.68
2 y = 88.68 -77.7
2 y = 10.98
y = 5.49
shirt ----$12.95
Tie----$5.49
m.ananth@hotmail.ca